PLEASE HELP ASAP!
Solve the problem by using a system of equations.
The income from a student production was $13,500. The price of a student ticket was $3, and nonstudent tickets were sold at $6 each. Four thousand tickets were sold. How many tickets of each kind were sold?
student tickets: ?
nonstudent tickets: ?

Answer by mananth(16946)

Student tickets x
Non student tickets y

x+y=4000.............1
3x+6y= 13500 .............2

multiply (1)by -6
Multiply (2) by 1
-6x-6y=-24000
3x+6y=13500
Add the two equations
-3x=-10500
/-3
x=3500
plug value of x in (1)
x+y=4000
3500+y= 4000
y=4000-3500
y=500
y = 500.00